OpenStudy (anonymous):

Given the function g(x) = 8x − 2, find g(−2).

OpenStudy (anonymous):

Basically all you have to do is substitute x for -2, so: g(-2) = 8*(-2) - 2 = -16 - 2 = -18

OpenStudy (anonymous):

oh, i knew that haha

OpenStudy (anonymous):

this is just setting x=-2

OpenStudy (anonymous):

i just realized that